This is an example ot the use of the subeqations package.

a^{2} + b^{2} = c^{2}  (1)

Now we start sub-numbering.

d^{2} + e^{2} = f^{2}  (2a)

We can refer to equation 1, 2 and 2a.

g^{2} + h^{2} = i^{2}  (2b)

This was equation 2b.

x = y + z  (2c)
u = v + w  (2d)

This was expression 2c, consisting of parts 2c and 2d. Now lets start a subeqnarray environment.

x = y + z  (3a)
u = v + w  (3b)

This was equation 3, with parts 3a and 3b.